OpenStudy (anonymous):

The small pipe has a radius of 18 feet and a height of 21 feet while the large pipe has a height of 28 feet. What is the volume of the large pipe in terms of π?

OpenStudy (ash2326):

@MykalStudy12 Is there a diagram?

OpenStudy (anonymous):

5,376π ft3 6,804π ft3 16,128π ft3 18,432π ft3

OpenStudy (campbell_st):

the ratio of the linear measurement is 21:28 or 3:4 or 1:(4/3) the volume of the small pipe is \[V = \pi \times18^2 \times 28\] \[V = 6804 \pi\] for linear ratios and volume, raise each term to the power of 3 \[(1)^3: (\frac{4}{3})^3 = 1: \frac{64}{27}\] so the volume of the larger pipe is \[V = 6804 \pi \times \frac{64}{27}\] I'll let you evalute it. the alternate method if to just use the linear ratio 3:4 to find the radius of the larger pipe... 18: x = 3:4